OpenStudy (anonymous):

Group Theory Could someone please explain to me why S3 is isomorphic to D6, and not to D3? Thanks :)

OpenStudy (anonymous):

no no one can explain it i hope

OpenStudy (anonymous):

at least in my understanding \(D_6\) has order 12 but maybe i am mistaken

OpenStudy (anonymous):

how exactly are you representing your elements of \(S_3\) ?

OpenStudy (anonymous):

whew i got scared for a moment

OpenStudy (nincompoop):

that may give an insight

OpenStudy (anonymous):

you can exhibit the isomorphism directly, but it depends on what you call your elements saying they are isomorphic amounts to saying it makes no difference, but you have to call them something

OpenStudy (nincompoop):

always a pleasure getting an insight from you, @satellite73

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

Sorry I'm confused, are you saying S3 isn't isomorphic to D6?

OpenStudy (anonymous):

Btw this isn't homework or anything, I'm just trying to better my understanding of isomorphisms, for example, I'm still unclear as to why Z(2)xZ(2) is isomorphic to V4? I understand the definition, or I should rather say, I have the definition, but I am finding it difficult to see why they are isomorphic, is it to do with no of elements? Thanks so much for any help :)

terenzreignz (terenzreignz):

What's V4?

terenzreignz (terenzreignz):

nvm, google :)

OpenStudy (nincompoop):

from what I've gathered, yes but I think the key is what satellite said

terenzreignz (terenzreignz):

They're isomorphic because they each have three elements of order 2? (About \(\large \mathbb{Z}_2\times \mathbb{Z}_2 \) and \(\large V_4\) )

terenzreignz (terenzreignz):

It seems odd because I've always defined the Klein-4 as the \(\large \mathbb{Z}_2 \times \mathbb{Z}_2\) itself, though I suppose it could go like this \[\Large V_4 = \left<a,b\right>\]where\[\Large a^2 = b^2 = (ab)^2 = 1\] Then just directly establish an isomorphism \[\large (0,0) \rightarrow1\\\large (1,0)\rightarrow a\\\large (0,1)\rightarrow b\\\large(1,1)\rightarrow ab\]

OpenStudy (anonymous):

Oh okay, that makes sense! Thankyou for the help guys :)

OpenStudy (nincompoop):

"an isomorphism f gives a one-to-one correspondence between the elements of two groups, which preserves the group operation. In other words, the fact that two groups are isomorphic, means that we can simply rename the elements of the first group (calling x by f(x)) and obtain exactly the second group. Obviously, the way we name elements is not important, so from the point of view of the group theory, isomorphic groups are identical."

OpenStudy (anonymous):

So because both S3 and D3 have the same elements, they are isomorphic? Because you can map any element of S3 in D3..is that right?

OpenStudy (nincompoop):

injective mapping, yes

terenzreignz (terenzreignz):

bijective?

OpenStudy (nincompoop):

mistyped (misspoke) bijective, yes

terenzreignz (terenzreignz):

The dihedral and permutation groups make me dizzy -.-

OpenStudy (nincompoop):

you got that right. it's like one of the things I have to read and re-read

OpenStudy (nincompoop):

why can't we just go back to the old days when we only bartered LMAO

OpenStudy (anonymous):

Haha nin ^^ so true

OpenStudy (nincompoop):

@KingGeorge

OpenStudy (kinggeorge):

Do you still need any help here?

OpenStudy (anonymous):

The only thing i'm struggling with now is what the differences are between the groups... Like D/S/Z/V?

OpenStudy (kinggeorge):

Well, \(D_{n}\) or \(D_{2n}\) depending on your notation, is the dihedral group, and basically corresponds to rotations/reflections of the \(n/2\)-gon or \(n\)-gon (depending on notation). \(S_n\) is the symmetric group, and is usually thought of as permutations of the set \(\{1,2,...,n\}\). This group has order \(n!\) and is very nice group to look for examples in. There is in fact a theorem that states: If a finite group has order \(n\), then it is isomorphic to a subgroup of \(S_{n!}\). \(\mathbb{Z}_n\) is simply the cyclic group of order \(n\). Finally, \(V_4\) is the Klein group. I haven't worked with this as much so I'm not an expert on it. However, it is normal in

OpenStudy (kinggeorge):

Whoops. That last sentence petered out there. It's normal in \(A_4\), which is itself normal in \(S_4\). \(A_n\) is the alternating group, and is a subgroup of \(S_n\). It's the set of even permutations, and is generated by the 3-cycles in \(S_n\). (If you haven't learned about normality or \(A_4\) yet, ignore this post